OpenStudy (anonymous):

Help me start this out please. I need to find the measure of x. Will medal as I always do :)

OpenStudy (anonymous):

it is a right triangle, such as this one. On side B is 12. On side C is 12. X is between the two.

OpenStudy (anonymous):

On side c is 13 My bad

OpenStudy (anonymous):

@Luigi0210 May you be able to help me? :)

OpenStudy (luigi0210):

So it's this? 

OpenStudy (anonymous):

Yes!! Without the a

OpenStudy (luigi0210):

Okayyyy, if you know your triangles, you should know that 5-12-13 is a special triangle like 3-4-5. So we can just replace a with 5. In order to find x, we could do either the sine, cosine, or tangent, your choice since we have all the sides.

OpenStudy (anonymous):

Im going to go with tan

OpenStudy (luigi0210):

Alright! So tangent is: \(\large \tan\theta=\frac{opposite}{adjacent} \) So for your triangle:  So \(\large \tan x=\frac{5}{12} \) Make sense?